OpenStudy (anonymous):

Find the slope of the line tangent to the graph at the given point. y= 2/4+x, x=8 What formula do I use to solve this one?

OpenStudy (anonymous):

this is the same as the last one. Just take the derivative first. then plug in 8 for x to get a slope

OpenStudy (anonymous):

d(u/v)=u'v-v'u/v^2 ? then substitute for u', v, v'?

OpenStudy (tkhunny):

\(\dfrac{d}{dx}(2/4+x) = \dfrac{d}{dx}(1/2 + x) = 1\) The slope of ALL tangent lines is 1. If you meant some other function, you will have to write it correctly.

OpenStudy (anonymous):

The answer is -1/72. I'm just not sure how to get to it.

OpenStudy (anonymous):

it cant be if the equation is what you provided

OpenStudy (tkhunny):

First, write it correctly. You need y = 2/(4+x). The parentheses are NOT optional.

OpenStudy (anonymous):

good catch @tkhunny

OpenStudy (anonymous):

from there y = 2(4+x)^-1 find the derivative of this

OpenStudy (anonymous):

can you do that

OpenStudy (anonymous):

Not really. The -1 is confusing me.

OpenStudy (anonymous):

its just a power, so use the power rule. y' = 2(-1)(x+4)^(-1-1)

OpenStudy (anonymous):

Ah, okay. I remember now.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

no problem